Determine the availability for each of these cases:

a. MTBF = 50 days, average repair time = 8 days. (Round your answer to 2 decimal places.)

Availability _______   

b. MTBF = 280 hours, average repair time = 10 hours. (Round your answer to 2 decimal places.)

Availability ________   
Answer #1

Availability = MTBF / (MTBF + MTTR)

a. Availability = 50 / 50+8 = 50/58 = 0.86

b. Availability = 280/280+10 = 280/290 = 0.96

PLEASE PLEASE PLEASE LEAVE A LIKE IF YOU FIND THIS HELPFUL. THANKS A LOT :))